# Importance sampling for Bayesian inference: polynomial-dimension dependent error bounds

### 02 Apr 2026 v1

@@ G.1 Mahalanobis-distance geometry
Throughout this section, geometric arguments are conducted w.r.t. the Mahalanobis norm $\|\cdot\|_{R}$. We assume the observation map $h\mathrel{\mathop{\ordinarycolon}}{\mathcal{X}}\subseteq\mathbb{R}^{d_{x}}\to\mathcal{Y}$ is uniformly bounded in the Euclidean norm; specifically, there exists a $d_{x}$-dependent constant
$M(d_{x})\mathrel{\mathop{\ordinarycolon}}=\sup_{x\in{\mathcal{X}}}\|h(x)\|<\infty.$
@@
$\|h(x)\|_{R}=\|R^{1/2}h(x)\|\leq\sqrt{\lambda_{1}(R)}\,\|h(x)\|\leq\sqrt{\lambda_{1}(R)}M(d_{x}).$
@@
$M_{R}^{d_{x}}\mathrel{\mathop{\ordinarycolon}}=\sup_{x\in{\mathcal{X}}}\|h(x)\|_{R}\leq\sqrt{\lambda_{1}(R)}M(d_{x}).$
@@
Applying the change of variables $\tilde{y}=y-h(x)$ to the inner integral, we obtain
$I(x)\mathrel{\mathop{\ordinarycolon}}=\int_{\mathcal{Y}}\frac{\phi^{2}\big(\psi(\|\tilde{y}\|_{R})\big)}{\phi\big(\psi(\|\tilde{y}+h(x)\|_{R}+M_{R}^{d_{x}})\big)}\,\mathrm{d}\tilde{y}$. (75)
@@
$\frac{\phi^{2}\big(\psi(\|\tilde{y}\|_{R})\big)}{\phi\big(\psi(\|\tilde{y}+h(x)\|_{R}+M_{R}^{d_{x}})\big)}\leq\frac{\phi^{2}\big(\psi(\|\tilde{y}\|_{R})\big)}{\phi\big(\psi(\|\tilde{y}\|_{R}+2M_{R}^{d_{x}})\big)}.$ (76)
@@
$\mathbb{E}\left[\|\ell_{Y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{2}\right]\leq\int_{{\mathcal{X}}}\left(\int_{\mathcal{Y}}k(y)\,\mathrm{d}y\right)\pi_{0}(\mathrm{d}x).$ (78)
@@
$\mathbb{E}\left[\|\ell_{Y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{2}\right]\leq\int_{\mathcal{Y}}k(y)\,\mathrm{d}y.$ (79)
@@
G.3.1 Case 1: Exponential radial profiles
@@
$\mathcal{J}\mathrel{\mathop{\ordinarycolon}}=\mathrm{C}\,\mathcal{S}_{R}^{d_{y}}\int_{0}^{\infty}\exp\left(\psi(r+2M_{R}^{d_{x}})-2\psi(r)\right)r^{d_{y}-1}\,\mathrm{d}r.$ (81)
The integrability is determined by the asymptotic behavior of the exponent $Q(r)\mathrel{\mathop{\ordinarycolon}}=a[(r+2M_{R}^{d_{x}})^{\beta}-2r^{\beta}]$. Factoring out the dominant term, we obtain
@@
    $\mathcal{J}_{1}\leq\mathrm{C}\,\mathcal{S}_{R}^{d_{y}}\exp\left(a(6M_{R}^{d_{x}})^{\beta}\right)\frac{(4M_{R}^{d_{x}})^{d_{y}}}{d_{y}}.$ (82)
@@
    $\mathcal{J}_{2}\leq\frac{\mathrm{C}\,\mathcal{S}_{R}^{d_{y}}}{\beta b^{d_{y}/\beta}}\Gamma\left(\frac{d_{y}}{\beta},b(4M_{R}^{d_{x}})^{\beta}\right),$ (83)
@@
G.3.2 Case 2: Polynomial-tail profiles
@@
$\mathbb{E}\left[\|\ell_{Y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{2}\right]\leq 2^{\alpha}(1+a(2M_{R}^{d_{x}})^{2})^{\alpha}=\mathrel{\mathop{\ordinarycolon}}K_{2}^{d_{x}},$ (85)
@@
FG and JM acknowledge the partial support of the Office of Naval Research (award N00014-22-1-2647), Spain’s Agencia Estatal de Investigación (award PID2021-125159NB-I00) and Comunidad de Madrid (IDEA-CM project, ref. TEC-2024/COM-89). The work of VE is supported by ARL/ARO under grant W911NF-22-1-0235 and by the Advanced Research + Invention Agency (ARIA).

### 28 May 2026 v2

@@ H.1 Mahalanobis-distance geometry
Throughout this section, geometric arguments are conducted w.r.t. the Mahalanobis norm $\|\cdot\|_{R}$. We assume the observation map $h^{d_{x}}\mathrel{\mathop{\ordinarycolon}}{\mathcal{X}}\subseteq\mathbb{R}^{d_{x}}\to\mathcal{Y}$ is uniformly bounded in the Euclidean norm; specifically, there exists a $d_{x}$-dependent constant
$M(d_{x})\mathrel{\mathop{\ordinarycolon}}=\sup_{x\in{\mathcal{X}}}\|h^{d_{x}}(x)\|<\infty.$
@@
$\|h^{d_{x}}(x)\|_{R}=\|R^{1/2}h^{d_{x}}(x)\|\leq\sqrt{\lambda_{1}(R)}\,\|h^{d_{x}}(x)\|\leq\sqrt{\lambda_{1}(R)}M(d_{x}).$
@@
$M_{R}^{d_{x}}\mathrel{\mathop{\ordinarycolon}}=\sup_{x\in{\mathcal{X}}}\|h^{d_{x}}(x)\|_{R}\leq\sqrt{\lambda_{1}(R)}M(d_{x}).$
@@
Applying the change of variables $\tilde{y}=y-h^{d_{x}}(x)$ to the inner integral, we obtain
$I(x)\mathrel{\mathop{\ordinarycolon}}=\int_{\mathcal{Y}}\frac{\phi^{2}\big(\psi(\|\tilde{y}\|_{R})\big)}{\phi\big(\psi(\|\tilde{y}+h^{d_{x}}(x)\|_{R}+M_{R}^{d_{x}})\big)}\,\mathrm{d}\tilde{y}$.
@@
$\frac{\phi^{2}\big(\psi(\|\tilde{y}\|_{R})\big)}{\phi\big(\psi(\|\tilde{y}+h^{d_{x}}(x)\|_{R}+M_{R}^{d_{x}})\big)}\leq\frac{\phi^{2}\big(\psi(\|\tilde{y}\|_{R})\big)}{\phi\big(\psi(\|\tilde{y}\|_{R}+2M_{R}^{d_{x}})\big)}.$
@@
$\mathbb{E}\left[\|\ell_{Y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{2}\right]\leq\int_{{\mathcal{X}}}\left(\int_{\mathcal{Y}}k(y)\,\mathrm{d}y\right)\pi_{0}(\mathrm{d}x).$
@@
$\mathbb{E}\left[\|\ell_{Y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{2}\right]\leq\int_{\mathcal{Y}}k(y)\,\mathrm{d}y.$
@@
H.3.1 Case 1: Exponential radial profiles
@@
$\mathcal{J}\mathrel{\mathop{\ordinarycolon}}=\mathrm{C}\,\mathcal{S}_{R}^{d_{y}}\int_{0}^{\infty}\exp\left(\psi(r+2M_{R}^{d_{x}})-2\psi(r)\right)r^{d_{y}-1}\,\mathrm{d}r.$
The integrability is determined by the asymptotic behavior of the exponent
$Q(r)\mathrel{\mathop{\ordinarycolon}}=a[(r+2M_{R}^{d_{x}})^{\beta}-2r^{\beta}].$
Factoring out the dominant term, we obtain
@@
    $\mathcal{J}_{1}\leq\mathrm{C}\,\mathcal{S}_{R}^{d_{y}}\exp\left(a(6M_{R}^{d_{x}})^{\beta}\right)\frac{(4M_{R}^{d_{x}})^{d_{y}}}{d_{y}}.$
@@
    $\mathcal{J}_{2}\leq\frac{\mathrm{C}\,\mathcal{S}_{R}^{d_{y}}}{\beta b^{d_{y}/\beta}}\Gamma\left(\frac{d_{y}}{\beta},b(4M_{R}^{d_{x}})^{\beta}\right),$
@@
H.3.2 Case 2: Polynomial-tail profiles
@@
$\mathbb{E}\left[\|\ell_{Y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{2}\right]=K_{2}^{d_{x}}\leq 2^{\alpha}(1+a(2M_{R}^{d_{x}})^{2})^{\alpha},$
@@
FG and JM acknowledge the support of Spain’s Agencia Estatal de Investigación (ref. PID2024-158181NB-I00 NISA), funded by MCIN/AEI/10.13039/501100011033 and by the ERDF (“A way of making Europe”); Comunidad de Madrid (IDEA-CM project, ref. TEC-2024/COM-89), and the Office of Naval Research (award N00014-22-1-2647). The work of VE is supported by ARL/ARO under grant W911NF-22-1-0235 and by the Advanced Research + Invention Agency (ARIA).